OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____. (13/34, -15/34) (3/22, -7/22) (13/22, 1/22) got c

OpenStudy (anonymous):

mult 1st by 4: 28x -12y= 16 mult 2nd by -3: -6x +12 = -3 combine: 22x = 13 x = 13/22 plug into one of the originals 2(13/22) - 4y = 1 13/11 -1 = 4y 13/11-11/11 = 4y 2/11 = 4y y = 1/22 yes, c is correct Good Job!